OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that contains the median to the base of an isosceles triangle with vertices A (2, 1), B(8, 4), and C(5, 7)

OpenStudy (anonymous):

Use the distance formula to figure out which sides are the congruent sides. \[d = \sqrt{(x_{1} - x_{2})^{2} + (y_{1} - y_{2})^{2}}\] AB = 3√5 AC = 3√5 BC = 3√2 BC is the non congruent side, so find the midpoint of that side. Midpoint of BC = (6.5, 5.5) Find the slope of the line 5.5 - 1/6.5 - 2 = 1 Plug it into the point-slope form. \[y - y_{1} = m(x - x_{1})\] y - 1 = 1(x - 2) y - 1 = x - 2 x - y = 1

OpenStudy (anonymous):

Thanks so much! :D

OpenStudy (anonymous):

You're welcome :)